In 2015, Juan won his second Indy 500, with a time of 3:05:57 for the 500 mi race. At one point, Juan was 20 mi closer to the finish than to the start. How far had Juan travel at that point?
Found 2 solutions by ikleyn, Alan3354:
Answer by ikleyn(52787)

.
x + y = 500  miles    (1)
x - y =  20  miles    (2)


Add the equations (both sides).


2x = 520  ====>  x = 520/2 = 260.


Answer.  260 miles.


Answer by Alan3354(69443)

In 2015, Juan won his second Indy 500, with a time of 3:05:57 for the 500 mi race. At one point, Juan was 20 mi closer to the finish than to the start. How far had Juan travel at that point?
-------------
The course is 2.5 miles long (around).
20 miles closer makes no sense.